OpenStudy (anonymous):

Solve each linear system in 2D (1) x+y=8 (2) x-y=13

jimthompson5910 (jim_thompson5910):

Hint: Add the equations to get x+y = 8 x-y = 13 ---------- 2x+0y = 21 2x = 21 x = ??? Now use this solution for x to get y

OpenStudy (anonymous):

Thanks !

jimthompson5910 (jim_thompson5910):

np